OpenStudy (anonymous):

use power reducing formula to write cos^4 x sin^4 x in terms of the first power of cosine

OpenStudy (anonymous):

Recall \[\cos^2(x) + \sin^2(x)=1\]

OpenStudy (anonymous):

\[\sin^2(x)=1-\cos^2(x)\]

OpenStudy (anonymous):

\[\cos^4(x)(1-\cos^2(x))^2\]

OpenStudy (anonymous):

\[=\cos^4(x)(1-2\cos^2(x)+\cos^4(x)\]

OpenStudy (anonymous):

or you may use \[\sin^2(x)=\frac{1-\cos(2x)}{2}\]

OpenStudy (anonymous):

this way you can reduce the power